I need help with this problem please:
1/2x-2/3= 4x

Answer by stanbon(75887)

1/2x-2/3= 4x
-----
Multiply thru by 6 to get:
3x - 4 = 24x
21x = -4
x = -4/21
==============
Cheers,
Stan H.